OpenStudy (lacypennelll):

Convert 9/20 to a decimal without dividing, do it step by step please to ^^

OpenStudy (mertsj):

To change to a decimal, the denominator must be a power of 10 since ours is a base 10 system. So do this: \[\frac{9}{20}=\frac{x}{100}\]

OpenStudy (lacypennelll):

9/20 = 45/100 = 0.45?

OpenStudy (mertsj):

Then change that number of hundredths to a decimal.

OpenStudy (mertsj):

Very good.

OpenStudy (lacypennelll):

Dang it there is that page glitch when it scrolls all the way down and then back up it's hard to fan like that DX